Tevin is 99 years younger than Rachael. 8 years ago, Rachael's age was 4 times Tevin's age. How old is Tevin now?

Answer by ikleyn(52788)

.

X    = Tevin now

x+99 = Rachael


8 years ago . . . 


    (x+99) -8 = 4*(x-8)


    x + 91 = 4x - 32

    91 + 32 = 4x - x

    123     = 3x

      x     = 123/3 = 41.      ANSWER


=============

By the way,  then Tevin age is  41+99 = 140.

Is Tevin an old tree ?   Or whale ?